OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = the quantity x minus seven divided by the quantity x plus three. and g(x) = quantity negative three x minus seven divided by quantity x minus one.

OpenStudy (anonymous):

\[f(x)=(\frac{ x-7 }{ x+3}) and g(x)=(\frac{ -3x-7 }{ x-1})\]

OpenStudy (anonymous):

put the g function in first then look at the f function where ever you see an x put in the whole of the g function it is untidy but simplifies to -x

OpenStudy (anonymous):

the first x in f(x) has to be replaced by {-3x-7}/(x-1) dont forget the -7 following

OpenStudy (johnweldon1993):

So yes, as @jamaica25 has said for \[\large f(g(x))\] you would take the g(x) function \[\large g(x) = \frac{-3x - 7}{x - 1}\] And plug that into your f(x) function wherever you see an 'x' So for this...we have \[\large f(x) = \frac{x - 7}{x + 3}\] which will now become \[\large f(g(x)) = \frac{\frac{-3x - 7}{x - 1} - 7}{\frac{-3x - 7}{x - 1} + 3}\] bad to look at now...but lets simplify it We need a common denominator on both the top and the bottom *which are both x - 1 so \[\large \frac{\frac{-3x - 7 - 7x + 7}{x - 1}}{\frac{-3x - 7 + 3x - 3}{x - 1}}\] Now those \(\large x - 1\) cancel and we have \[\large \frac{-3x - 7 - 7x + 7}{-3x - 7 + 3x - 3}\] Combine like terms \[\large \frac{-10x}{-10}\] which does indeed = x

OpenStudy (johnweldon1993):

And it would be the same process for the g(f(x)) one...so let me know if anything trips you up :)

OpenStudy (anonymous):

Thank you so much!!!!

OpenStudy (johnweldon1993):

Not a problem :)